### v1

# Efficient Reinforcement Finetuning via Adaptive Curriculum Learning

Abstract
Reinforcement finetuning (RFT) has shown great potential for enhancing the mathematical reasoning capabilities of large language models (LLMs), but it is often sample- and compute-inefficient, requiring extensive training. In this work, we introduce AdaRFT (Adaptive Curriculum Reinforcement Finetuning), a method that significantly improves both the efficiency and final accuracy of RFT through adaptive curriculum learning. AdaRFT dynamically adjusts the difficulty of training problems based on the model’s recent reward signals, ensuring that the model consistently trains on tasks that are challenging but solvable. This adaptive sampling strategy accelerates learning by maintaining an optimal difficulty range, avoiding wasted computation on problems that are too easy or too hard. AdaRFT requires only a lightweight extension to standard RFT algorithms like Proximal Policy Optimization (PPO), without modifying the reward function or model architecture. Experiments on competition-level math datasets—including AMC, AIME, and IMO-style problems—demonstrate that AdaRFT significantly improves both training efficiency and reasoning performance. We evaluate AdaRFT across multiple data distributions and model sizes, showing that it reduces the number of training steps by up to 2 and improves accuracy by a considerable margin, offering a more scalable and effective RFT framework.
1 Introduction
Reinforcement Finetuning (RFT) has emerged as a powerful technique for aligning large language models (LLMs) with task-specific goals, particularly in domains such as mathematics and code generation where correctness is well defined (DeepSeek-AI et al., 2025; OpenAI et al., 2024). By optimizing a policy model with reward signals that reflect task success, RFT enables more targeted learning than supervised finetuning (SFT) alone. However, despite its promise, RFT remains sample-inefficient and computationally expensive. Its training involves repeated rollout generation, reward computation, and policy updates—making it costly and difficult to scale (Ahmadian et al., 2024; Kazemnejad et al., 2024; Li et al., 2024; Hu, 2025; Cui et al., 2025). Recent efforts to address RFT inefficiency have focused on algorithmic simplification (e.g., RAFT (Dong et al., 2023), GRPO (DeepSeek-AI et al., 2025), ReMax (Li et al., 2024)), and data-centric strategies (e.g., LIMO (Ye et al., 2025), LIMR (Li et al., 2025)). While these approaches improve sample or compute efficiency, they often introduce trade-offs: algorithmic simplifications may increase variance or limit stability, and static data filtering or scoring can be brittle, computationally heavy, or model-specific. Moreover, most methods’ success relies on fixed datasets or training schedules, which can be suboptimal in non-uniform or imbalanced data regimes.
To overcome these limitations, we propose AdaRFT, a reinforcement finetuning method based on adaptive curriculum learning (Bengio et al., 2009), which dynamically adjusts training set difficulty to match the model’s evolving skill level. The algorithm is shown in Alg. 1. To illustrate the motivation, consider asking a grade school student to solve a problem from the International Mathematical Olympiad (IMO). Without the necessary background and tools, the student would struggle to engage meaningfully with the task, let alone learn from it. Inspired by this analogy, we hypothesize that LLMs trained via RFT may face a similar challenge: when presented with tasks that exceed their current capabilities, the learning signal becomes weak or ineffective. On the other hand, tasks that are too simple may not sufficiently challenge the model to improve. In both scenarios, a mismatch between task difficulty and model ability leads to inefficient learning.
AdaRFT addresses this challenge by dynamically selecting training problems that match the model’s current skill level. Each problem in the dataset is preassigned a difficulty score, which can be either estimated automatically or derived from human annotations. During training, we maintain a target difficulty level that is adjusted based on the model’s recent reward signal: as the model improves, the target difficulty increases; if performance drops, it decreases. At each step, the model is finetuned on problems whose difficulty is closest to this target, ensuring a steady progression through increasingly challenging but solvable tasks.
Unlike prior work that simplifies the RL algorithm or precomputes model-specific learning-impact metrics, AdaRFT adds only a lightweight curriculum scheduler on top of standard RFT algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017), with no modification to reward functions and no model-dependent preprocessing. While AdaRFT is compatible with a variety of policy optimization methods (e.g., PPO, GRPO), in this work, we instantiate it using PPO for all experiments to demonstrate its effectiveness. We denote this specific instantiation as AdaRFT (PPO). We evaluate AdaRFT on a dataset spanning a wide range of competition-level math problems, including AMC, AIME, and IMO-style questions. Across multiple training distributions and two model sizes, adaptive curriculum learning significantly improves both training efficiency and final performance. Gains are especially notable in imbalanced data regimes, where static sampling often fails. AdaRFT can speed up training by up to 2, offering a practical and scalable path to more efficient RFT in structured reasoning tasks.
2 Related Work
Efficient Reinforcement Finetuning.
Most RFT pipelines are based on Proximal Policy Optimization (PPO) (Schulman et al., 2017), with recent variants such as RAFT (Dong et al., 2023), GRPO (DeepSeek-AI et al., 2025), REINFORCE++ (Hu, 2025), and ReMax (Li et al., 2024) aiming to reduce computational cost by simplifying the underlying RL algorithm. These methods generally improve efficiency and stability, but often introduce trade-offs. For example, GRPO eliminates the value function by estimating advantages through relative comparisons within groups of outputs. While this removes the need for a critic and reduces complexity, it can lead to instability due to noisier advantage estimates, higher variance in updates, and increased sample requirements (DeepSeek-AI et al., 2025; Hu, 2025). In parallel, data-centric strategies have emerged as promising alternatives for efficient finetuning. Techniques like LIMO (Ye et al., 2025) and s1 (Muennighoff et al., 2025) demonstrate that small, carefully curated supervised finetuning datasets can achieve strong performance while using orders of magnitude less data. However, they rely heavily on effective data selection and prompt design, which may hinder generalization to new tasks. LIMR (Li et al., 2025) also aims to reduce RFT sample requirements by scoring training examples using a learning impact measurement (LIM). Although this allows subsequent training with fewer samples, computing LIM scores requires a full training run on the entire dataset. Moreover, despite using fewer samples afterwards, models still take roughly the same number of steps to converge, limiting the overall efficiency gains. LIMR is also model-dependent, meaning that LIM scores must be recomputed when switching to a different model, further increasing overhead. In contrast to these approaches, AdaRFT introduces an adaptive curriculum learning strategy that adaptively selects training problems based on the model’s current performance. Rather than simplifying the RL algorithm or relying on static data selection, we continuously adjust the training difficulty to match the model’s evolving capability. This curriculum-aware training leads to faster convergence, improved accuracy, and higher sample and compute efficiency, especially in non-uniform training distributions, while maintaining stability and generalization across tasks.
Curriculum Learning for RL.
Curriculum learning (CL) (Bengio et al., 2009) is a training strategy that structures learning by presenting training samples in a meaningful progression, typically from simpler to more difficult tasks. In RL, task-specific curricula sort tasks based on predefined or learned difficulty metrics to guide agents through progressively harder challenges (Zaremba & Sutskever, 2015; Justesen et al., 2018; Wang et al., 2019). Teacher-guided curricula treat task selection as a bandit or partially observable Markov decision process problem (POMDP), where a teacher agent adaptively assigns tasks that maximize the student’s learning progress (Matiisen et al., 2017; Portelas et al., 2019). Self-play methods generate an automatic curriculum by having agents propose and solve tasks for each other, dynamically increasing task difficulty through competition (Sukhbaatar et al., 2018). Goal generation techniques, such as Goal GANs, produce goals of intermediate difficulty based on current policy performance to optimize learning in sparse-reward settings (Florensa et al., 2018). Skill-based curricula organize unsupervised trajectories into a latent skill space to create transferable training tasks for meta-RL (Jabri et al., 2019). Distillation-based approaches, like progressive networks (Rusu et al., 2022) and Mix & Match (Czarnecki et al., 2018), transfer knowledge from simpler agents to more complex ones through shared architectures and imitation losses. These diverse approaches demonstrate the value of structured task progression in improving generalization and sample efficiency across complex RL domains. While curriculum learning has been widely studied in classical RL, its application to RFT of LLMs remains largely unexplored. Our work is among the first to introduce adaptive curriculum learning in the RFT setting, adapting problem difficulty in real-time based on model performance. This approach improves both sample efficiency and final performance, demonstrating that curriculum learning is a powerful tool for scaling RFT on complex reasoning tasks.
3 AdaRFT
We aim to improve the performance of a policy model for solving mathematical problems through adaptive curriculum learning. Fine-tuning on problems that are too easy or too hard leads to poor learning outcomes. Instead, the model should be trained on problems whose difficulty is close to the model’s current capability. We frame this as an adaptive curriculum learning problem and propose AdaRFT, which adaptively adjusts the target difficulty to keep training problems within a suitable difficulty range. AdaRFT is compatible with a variety of RL algorithms (e.g, GRPO, PPO); in this work, we instantiate it with PPO and refer to this variant as AdaRFT (PPO).
Let be a dataset of mathematical problems, each annotated with a precomputed difficulty score . For example, the MATH dataset (Hendrycks et al., 2021) includes preassigned difficulty levels ranging from 1 to 5. These difficulty scores can also be derived through other methods. One common approach is to use human expert annotations; for instance, problems from the International Mathematical Olympiad (IMO) are generally considered more difficult than those from the American Mathematics Competitions (AMC). Alternatively, difficulty can be estimated based on the model’s empirical success rate: problems that are frequently solved correctly receive lower difficulty scores, while problems that are rarely solved correctly are given higher scores. Another approach is to train a separate model, such as a regression or classification model, to estimate difficulty from features of the problem. Given a predefined difficulty for each problem, the objective is to train a policy that improves its problem-solving ability by dynamically adjusting the training curriculum according to the model’s current performance. Our proposed algorithm, AdaRFT, is shown in Algorithm 1.
3.1 Dynamic Curriculum Sampling
To construct an adaptive curriculum, we define a target difficulty , which represents the current target difficulty level for training. AdaRFT dynamically adjusts based on the model’s reward signal to maintain an optimal difficulty level for learning.
At each training step (see Algorithm 1), the algorithm computes the absolute difference between the target difficulty and the difficulty of each problem in the dataset (Step 1.4): for all . The batch of training problems is formed by selecting the problems with the smallest values of (Step 1.5), producing a batch: .
This ensures that the selected problems are closest to the model’s current target difficulty, focusing the learning process on problems that are neither too easy nor too hard.
3.2 Policy Update
The selected batch is used to train the policy model , which generates responses (Step 1.6): . A reward signal is then computed based on the correctness of the model’s output (Step 1.7):
The average reward over the batch is computed as (Step 1.7): .
The policy can then be updated using a reinforcement learning algorithm such as PPO, GRPO, or REINFORCE++ (Step 1.8): .
3.3 Target Difficulty Update
To adapt the curriculum dynamically, the target difficulty is updated based on the average reward. If the model performs well on the current difficulty level (high reward), the target difficulty increases, making the training problems harder. Conversely, if the model performs poorly, the target difficulty decreases. This dynamic update mechanism lies at the core of AdaRFT’s curriculum adaptation strategy.
The update rule (Step 1.9) is defined as:
Here, are hyperparameters: is the step size for adjusting the target difficulty, controls the sensitivity of the update, and is the target reward level, representing the desired success rate. The function ensures smooth updates and prevents large jumps in difficulty by saturating for large deviations, while the “clip” function constrains the target difficulty within the valid range . These bounds can be manually specified or automatically derived from the training set, for example by taking the minimum and maximum of the difficulty scores . Intuition and guidance for selecting these hyperparameters are discussed in Section 4.3.
4 Experiments
4.1 Difficulty Estimation
Accurate estimation of problem difficulty is critical for AdaRFT. For difficulty estimation, we use the Qwen 2.5 MATH 7B model (Qwen et al., 2025).111https://huggingface.co/Qwen/Qwen2.5-Math-7B A model that is too strong (e.g., OpenAI o1 (OpenAI et al., 2024), DeepSeek R1 (DeepSeek-AI et al., 2025)) would solve most problems on the first attempt, leading to poor discrimination between easy and hard problems. Conversely, a model that is too weak (e.g., LLaMA 3.3 1B (Grattafiori et al., 2024)) would fail to solve most problems even after multiple attempts, limiting the signal required for curriculum adaptation.
We select Qwen 2.5 MATH 7B as the estimator because it demonstrates a balanced solving ability — it can solve moderately difficult problems while still failing on more complex problems, thereby providing a more informative difficulty signal. For each problem, the difficulty score is computed as:
where is the number of attempts allowed for the model. In our setup, we use . This approach ensures that difficulty scores reflect the true problem-solving capability of a moderately capable model, preventing over- or under-estimation of difficulty.
To evaluate the stability of our difficulty estimation process, we simulate how confidence varies with different numbers of samples. For each problem, we treat the full set of 128 rollouts as the ground-truth difficulty estimate and compute how often sub-sampled estimates fall within a tolerance of . Specifically, we run 10 random sampling trials per sample size and average the confidence across all problems in the dataset. As shown in Figure 1(a), even with as few as 64 samples, the estimated difficulty remains within of the full estimate over 90% of the time. With just 40 samples, the confidence remains around 80%. These results indicate that accurate and robust difficulty estimation can be achieved with significantly fewer rollouts, reducing the computational burden of large-scale curriculum construction.
To further validate the reliability of our difficulty estimates, we examined their alignment with the human-labeled difficulty levels provided in the MATH dataset. The MATH dataset comprises 12,500 competition-level mathematics problems sourced from contests such as the AMC 10, AMC 12, and AIME. Each problem is categorized into one of five difficulty levels, following the classification system used by the Art of Problem Solving (AoPS) community.222https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki%3ACompetition_ratings In this dataset, a subject’s easiest problems are assigned a difficulty level of 1, while the hardest problems are assigned a difficulty level of 5. For instance, the initial problems in an AMC 8 exam are often labeled as level 1, whereas AIME problems are typically labeled as level 5. As shown in Figure 1(b), there is a clear downward trend in average solved percentage as the labeled difficulty level increases, ranging from 86.0% at level 1 to 52.7% at level 5. This strong negative correlation indicates that the model’s empirical performance is consistent with the intended difficulty stratification, reinforcing the usefulness of both the labeled difficulty levels and our estimation approach for guiding curriculum learning.
4.2 Dataset
We use the DeepScaleR dataset (Luo et al., 2025) as the training set. DeepScaleR compiles problems from multiple sources, including the American Invitational Mathematics Examination (AIME) from 1984 to 2023 and the American Mathematics Competitions (AMC) prior to 2023. In addition, the dataset includes problems from the Omni-MATH (Gao et al., 2024) and Still datasets (Team, 2025), which feature problems from various national and international math competitions. This results in a diverse and challenging training set, covering a wide range of mathematical domains and difficulty levels.
In practice, we do not have control over the exact difficulty distribution of the data collected for training. This motivates our investigation into how different difficulty distributions influence AdaRFT. To this end, we construct three distinct distributions from the DeepScaleR dataset. The first is a skew-difficult distribution, where most problems are challenging. The second is a skew-easy distribution, where most problems are relatively easy. The third is a uniform distribution, where problems are evenly balanced across all difficulty levels, ensuring a consistent representation of easy, moderate, and hard problems. Each of these three distributions includes 10,000 samples. The data distribution for each setting is shown in Figure 2.
For evaluation, we use six benchmark datasets to assess the model’s performance across different levels of difficulty and mathematical reasoning. The first benchmark, MATH 500 (Lightman et al., 2023), is a subset of the MATH dataset (Hendrycks et al., 2021) containing 500 representative problems designed to test a model’s general mathematical capability. GSM8K (Cobbe et al., 2021) is a set of grade-school math problems. OlympiadBench (He et al., 2024) includes a collection of problems from Olympiad-level mathematics and physics competitions. Minerva Math (Lewkowycz et al., 2022) is a curated set of undergraduate-level math problems that assess complex mathematical reasoning and symbolic manipulation. AMC 23 and AIME 24 include problems from the 2023 American Mathematics Competitions and the 2024 American Invitational Mathematics Examination, respectively. Since AMC 23 contains only 40 problems and AIME 24 only 30, we report accuracy as the average over 8 sampled responses per problem to ensure stable estimates. Together, these datasets span elementary, high school, and advanced competition-level math, providing a comprehensive evaluation of the model’s reasoning abilities.
4.3 Training Setup
We trained two models on the three difficulty-based distributions of the DeepScaleR dataset described in Section 4.2: Qwen 2.5 7B333https://huggingface.co/Qwen/Qwen2.5-7B and Qwen 2.5 MATH 1.5B444https://huggingface.co/Qwen/Qwen2.5-Math-1.5B. This setup allows us to evaluate the effectiveness of AdaRFT on models with different initial performance levels when exposed to skew-hard, skew-easy, and uniform problem distributions. Both models were trained using two approaches: the standard PPO algorithm and AdaRFT (PPO), our method that integrates adaptive curriculum learning with PPO (see Section 3). For AdaRFT, the target difficulty was dynamically adjusted throughout training based on the model’s reward signal. This adjustment ensured that the model consistently encountered problems that matched its current skill level, preventing the model from being overwhelmed by overly difficult problems or stagnating on problems that were too easy.
The training batch size was set to , with the target reward set to 0.5 to promote learning at a balanced success rate. The sensitivity parameter and step size were tuned using a validation set to ensure stable curriculum updates. We set , , and the initial target difficulty . The step size acts as a scaling factor between the reward signal and the difficulty metric. Since the difficulty metric ranges from 0 to 100 and the reward ranges from 0 to 1, a target reward implies that the maximum reasonable adjustment to the difficulty metric should be around 50. Therefore, we set to scale the reward signal appropriately to the difficulty range. The sensitivity parameter controls the slope of the function. Setting to 2 makes the function behave approximately linearly when the difference between the average reward and the target reward is small. The intuition behind using the function is that when the average reward is close to the target reward, a roughly linear adjustment is appropriate. However, when the average reward deviates significantly from the target reward, linear adjustments may be too large, leading to instability. The function smooths out these adjustments, allowing for more controlled changes when the difference is large while maintaining sensitivity when the difference is small. Both models were trained on 8 A100 GPUs for approximately 100 steps. The implementation details can be found in Appendix A.
5 Result
We evaluate the performance of standard PPO and AdaRFT (PPO) across multiple training setups and two model sizes: Qwen 2.5 MATH 1.5B and Qwen 2.5 7B. Figure 3 presents the learning curves averaged across six benchmarks, while Table 1 and 2 provide a detailed breakdown of accuracy and training efficiency at step 60 and step 100. On average, models trained with AdaRFT (PPO) outperform their PPO-only counterparts in both final accuracy and training efficiency. This improvement is particularly notable in non-uniform data distributions, where curriculum adaptation is most beneficial.
5.1 Training Efficiency
As shown in Figure 3 and Table 1, models with AdaRFT (PPO) require substantially fewer training steps to reach the same level of performance as standard PPO. Specifically, we measure how many additional training steps PPO requires to match the performance that AdaRFT (PPO) attains at step 60. Because evaluation is performed only every 5 steps during training, we estimate these step savings by interpolating the smoothed accuracy curves. We apply exponential smoothing with a factor of 0.7 to reduce evaluation noise and obtain stable estimates of performance over time.
For Qwen 2.5 Math 1.5B, PPO requires 26 additional steps in the skew-difficult setting and 29 in the uniform setting to reach the performance achieved by AdaRFT (PPO) at step 60—corresponding to 43.7% and 48.3% reductions in training steps, respectively. Even in the skew-easy setting, PPO lags behind by 9 steps. The efficiency gains are even more pronounced with the larger Qwen 2.5 7B model: in the skew-difficult setting, PPO requires 44 additional steps, a 73.2% increase, and in the skew-easy setting, 64 additional steps, a 106.5% increase. On uniformly sampled data, both methods converge at a similar rate, though AdaRFT (PPO) maintains a slight advantage.
Notably, these gains are accompanied by a reduction in computational cost per step. As shown in Table 1, AdaRFT (PPO) consistently achieves lower average time per training step than standard PPO across all settings. While the underlying causes are discussed in Section 6, this result highlights that AdaRFT improves not only convergence speed but also overall training throughput, making it both sample- and compute-efficient.
| Model | Setup | Method |
| Qwen 2.5 Math 1.5B | skew-difficult | PPO | 69.14 | 60.20 | 20.36 | 12.13 | 45.31 | 7.50 | 35.77 | 26 (43.67%) | 134.44 | ||||||||||||||||
| AdaRFT | 71.42 | 63.00 | 20.51 | 15.07 | 45.62 | 10.83 | 37.74 | 124.66 | |||||||||||||||||||
| uniform | PPO | 70.74 | 61.20 | 19.61 | 14.34 | 45.00 | 11.67 | 37.09 | 29 (48.32%) | 129.24 | |||||||||||||||||
| AdaRFT | 72.55 | 65.20 | 21.10 | 13.24 | 47.19 | 14.17 | 38.91 | 120.21 | |||||||||||||||||||
| skew-easy | PPO | 70.58 | 61.40 | 20.95 | 14.34 | 47.50 | 12.92 | 37.95 | 9 (14.62%) | 119.99 | |||||||||||||||||
| AdaRFT | 72.25 | 63.60 | 19.76 | 15.81 | 46.88 | 14.17 | 38.74 | 120.91 | |||||||||||||||||||
| Qwen 2.5 7B | skew-difficult | PPO | 88.17 | 69.40 | 21.99 | 25.00 | 45.63 | 9.58 | 43.29 | 44 (73.18%) | 258.87 | ||||||||||||||||
| AdaRFT | 89.08 | 72.20 | 24.52 | 24.26 | 48.12 | 13.33 | 45.25 | 247.73 | |||||||||||||||||||
| uniform | PPO | 89.61 | 71.00 | 23.92 | 23.53 | 49.06 | 9.17 | 44.38 | 3 (5.27%) | 254.38 | |||||||||||||||||
| AdaRFT | 88.32 | 73.40 | 24.81 | 22.06 | 45.63 | 15.83 | 45.01 | 241.64 | |||||||||||||||||||
| skew-easy | PPO | 88.86 | 72.80 | 24.07 | 22.79 | 48.13 | 13.33 | 45.00 | 64 (106.5%) | 250.22 | |||||||||||||||||
| AdaRFT | 88.48 | 73.40 | 23.77 | 23.53 | 49.69 | 12.92 | 45.30 | 253.21 |
5.2 Model Performance
In addition to improving efficiency, AdaRFT (PPO) also improves the final model performance. As shown in Table 2, at the end of training (step 100), curriculum learning yields consistent improvements in final accuracy across all configurations. As shown in Table 2, AdaRFT (PPO) outperforms standard PPO in nearly every setting, with especially strong gains on the Qwen 2.5 Math 1.5B model. The reported averages reflect accuracy across six diverse benchmarks: GSM8K, MATH 500, OlympiadBench, Minerva Math, AMC 23, and AIME 24.
| Model | Setup | Method |
Average | |||||||||||
| Qwen 2.5 Math 1.5B | skew-difficult | PPO | 69.67 | 64.60 | 20.65 | 12.87 | 47.50 | 9.17 | 37.41 | |||||||||||
| AdaRFT | 74.00 | 66.40 | 20.36 | 15.07 | 55.00 | 12.08 | 40.48 | |||||||||||||
| uniform | PPO | 71.95 | 65.20 | 21.10 | 15.81 | 42.50 | 6.67 | 37.20 | ||||||||||||
| AdaRFT | 74.53 | 66.20 | 21.99 | 14.34 | 57.50 | 12.08 | 41.11 | |||||||||||||
| skew-easy | PPO | 72.71 | 67.40 | 19.17 | 13.97 | 45.00 | 12.50 | 38.46 | ||||||||||||
| AdaRFT | 73.62 | 66.20 | 19.91 | 13.97 | 55.00 | 9.17 | 39.18 | |||||||||||||
| Qwen 2.5 7B | skew-difficult | PPO | 89.69 | 71.20 | 23.33 | 23.53 | 50.00 | 11.25 | 44.17 | |||||||||||
| AdaRFT | 90.98 | 71.40 | 25.85 | 22.43 | 52.50 | 15.83 | 46.83 | |||||||||||||
| uniform | PPO | 89.31 | 72.40 | 23.63 | 25.37 | 42.50 | 15.00 | 44.70 | ||||||||||||
| AdaRFT | 90.14 | 72.60 | 24.96 | 24.26 | 55.00 | 14.58 | 46.92 | |||||||||||||
| skew-easy | PPO | 89.39 | 73.60 | 23.33 | 24.26 | 47.50 | 13.33 | 45.07 | ||||||||||||
| AdaRFT | 90.14 | 72.60 | 25.56 | 23.16 | 50.00 | 14.17 | 45.94 |
On skew-difficult data, the Qwen 2.5 Math 1.5B model improves from 37.41% (PPO) to 40.48% (AdaRFT (PPO)), a gain of over 3 percentage points in average accuracy. Similar improvements appear in the uniform setting, where AdaRFT (PPO) reaches 41.11%, compared to 37.20% with PPO. Even on skew-easy data, where the baseline performs well, curriculum learning still improves performance, reaching 39.18% versus 38.46%. For the larger Qwen 2.5 7B model, final accuracy gains are also consistent, though slightly more modest. In the skew-difficult setting, AdaRFT (PPO) improves from 44.17% to 46.83%. In the uniform setting, accuracy rises from 44.70% to 46.92%, and in the skew-easy case, from 45.07% to 45.94%. These results show that curriculum learning is effective even for stronger models, enhancing both stability and peak performance.
6 Discussion
6.1 Analysis on Efficiency and Accuracy
Sample efficiency and final performance.
AdaRFT substantially improves both sample efficiency and final model performance, particularly in non-uniform data distributions. By dynamically adjusting task difficulty during training, AdaRFT enables models to allocate learning capacity more effectively, focusing on simpler tasks early on and gradually transitioning to harder ones. This leads to faster convergence and better use of training steps. The performance gains are especially pronounced in the Qwen 2.5 Math 1.5B model, where curriculum structure plays a larger role in guiding learning. For the larger Qwen 2.5 7B model, improvements are still evident, though more modest—likely because the model’s increased capacity allows it to handle diverse and imbalanced tasks without as much external scaffolding.
Reduced training time per step.
AdaRFT also results in slightly lower average training time per step. This is largely due to the fact that easier problems require fewer tokens to solve. For example, an arithmetic reasoning question from GSM8K might require only around 200 tokens for Qwen 2.5 7B to reach a correct answer, whereas a competition-level math problem from AIME could require around 2000 tokens—a 10 difference in rollout length. The total token length affects multiple components of the training step, including the rollout itself and the subsequent PPO update. While PPO update time does not scale linearly with sequence length due to batching and attention computation patterns, longer sequences still incur higher compute cost. As a result, curriculum learning’s tendency to prioritize shorter, easier problems early in training leads to shorter sequences on average, reducing per-step compute and improving overall training throughput.
When does curriculum learning help?
Our findings show that curriculum learning provides the greatest benefits under two key conditions: (1) imbalanced training distributions, and (2) limited model capacity. In skewed distributions, particularly the skew-difficulty settings, standard PPO often struggles to gain traction early in training due to insufficient reward signals. AdaRFT mitigates this by initially sampling easier problems, enabling the model to bootstrap capabilities before tackling harder content. For example, in the Qwen 2.5 Math 1.5B model trained on skew-difficult data, AdaRFT improves final accuracy by over 3 percentage points and reduces training steps by 44%. This suggests that AdaRFT is especially helpful when the model would otherwise be under-optimized due to overly difficult or unbalanced data.
Conversely, the benefits of AdaRFT are less pronounced when the model is strong enough or the data is already well-balanced. For instance, in the uniform setup with the Qwen 2.5 7B model, AdaRFT offers only modest improvements in convergence speed and final accuracy. Likewise, in the skew-easy setup for the smaller Qwen 2.5 Math 1.5B model, the curriculum yields relatively smaller gains. In both cases, the model is either already exposed to a representative distribution of task difficulties or finds most problems challenging enough, thus reducing the need for dynamic difficulty adjustment.
It is important to note that manual data curation and task scheduling, carefully designed for a specific model, could potentially achieve similar results by selecting a training sequence that aligns with the model’s learning capacity (Yu et al., 2025; Shen et al., 2025; Chen et al., 2025; Zeng et al., 2025). However, such methods require significant human effort, domain knowledge, and often need to be re-tuned for each new model or task. In contrast, AdaRFT requires no manual data curation or model-specific preprocessing. The curriculum automatically adapts to the model’s reward signal during training, making it broadly applicable across model scales and training distributions. This automatic adaptability not only saves engineering effort but also improves scalability and robustness in real-world training pipelines. Moreover, AdaRFT is particularly advantageous in fixed data settings, as it can tailor the difficulty schedule to match the capabilities of any model, whether strong or weak, without altering the dataset, providing a unified solution that generalizes across model architectures and skill levels.
6.2 Data Difficulty on Model Performance
To better understand the effect of data difficulty on model performance, we introduce two additional data distributions: easy-extreme and hard-extreme. Unlike the skew-difficult and skew-easy distributions, which still include a mix of difficulty levels, the easy-extreme and hard-extreme sets consist exclusively of the most polarized examples. Specifically, easy-extreme contains only the easiest samples with difficulty levels no greater than 15, while hard-extreme includes only the hardest samples with difficulty levels of at least 97. Each of these extreme distributions consists of approximately 8,000 samples, providing a focused and controlled evaluation of model behavior under minimal or maximal difficulty conditions. We trained a Qwen 2.5 7B model on each of the two extreme distributions using PPO, and compared their performance to models trained on the uniform distribution with PPO (Uniform) and with AdaRFT instantiated with PPO (Uniform + AdaRFT), as described in Section 5. The results are presented in Figure 4.
Accuracy.
The leftmost panel shows that uniform + AdaRFT achieves the highest overall accuracy throughout training, outperforming both uniform and the two extreme settings. This highlights the effectiveness of AdaRFT in guiding the model through an optimal difficulty progression. In contrast, hard-extreme struggles significantly, with a flat and lower trajectory, indicating that exposing the model only to very difficult problems limits learning progress. This suggests that without a gradual exposure strategy, models trained on only the hardest problems are unable to bootstrap their capabilities effectively.
Reward.
The reward trends provide important clues about learning dynamics. The easy-extreme setup achieves the fastest reward improvement during early training, surpassing both uniform and hard-extreme. In particular, easy-extreme consistently operates in a reward range between 0.4 and 0.6 during early training, which corresponds to a success rate that is both challenging and attainable. In contrast, the reward of the uniform and hard-extreme setup lingers below 0.2 in early training, leading to slower learning. This suggests that training on problems with intermediate difficulty—those that are neither trivially easy nor prohibitively hard—provides the most effective learning signal. Notably, AdaRFT is explicitly designed to exploit this insight: by setting the target reward , we encourage the model to train on problems that match this “productive struggle” zone. As shown by the uniform + AdaRFT curve, the algorithm successfully maintains an average reward near 0.5 throughout training, allowing the model to learn at an optimal pace. Notably, while the uniform setup eventually reaches a reward of nearly 0.5 by step 50, it does not result in faster learning. This is likely because the model is already fairly well trained by that stage, so the additional reward signal contributes less to further improvement. In contrast, the hard-extreme model receives almost no reward signal for most of the training, while the uniform setup shows slower and more gradual reward accumulation.
Response Length.
The response length panel reveals how the complexity of generated solutions evolves during training. The hard-extreme model consistently produces the longest responses, with length increasing steadily, reflecting the higher complexity and reasoning depth required by the hardest problems. In contrast, the easy-extreme setup maintains short and stable responses, consistent with its simpler problem set. The uniform and uniform + AdaRFT setups fall between these two extremes. Notably, uniform + AdaRFT shows a gradual increase in response length over time. This trend aligns with the behavior of the curriculum learning algorithm: as the model improves, it is exposed to increasingly difficult problems, which naturally demand more elaborate reasoning and longer solutions. This dynamic suggests that response length can serve as a useful proxy for problem difficulty and reasoning complexity during training.
Difficulty.
Finally, the difficulty panel illustrates how problem difficulty evolves under each setup. The easy-extreme and hard-extreme curves remain flat, confirming that these datasets contain only problems from the tail ends of the difficulty spectrum (i.e., and , respectively). The uniform curve is centered around 50, as expected, while uniform + AdaRFT shows a steady increase in difficulty over time. This adaptive progression confirms that curriculum learning effectively steers the model from easier to harder problems, aligning difficulty with the model’s evolving capabilities.
7 Conclusion
We propose AdaRFT, an adaptive curriculum learning strategy for reinforcement finetuning (RFT) that dynamically matches problem difficulty to a model’s evolving skill level. By adjusting a target difficulty based on reward feedback, AdaRFT improves both sample and compute efficiency without modifying the reward function or underlying RL algorithm. Experiments across multiple data regimes and model sizes show consistent gains in convergence speed and final accuracy, especially in imbalanced training distributions. This lightweight, scalable approach highlights the value of curriculum-aware training for efficient and robust alignment in structured reasoning tasks.
Acknowledgement
The authors would like to thank the members of the LIME lab and the USC NLP group for their helpful insights and feedback.
Future Works
Despite these encouraging results, there are important limitations. Our curriculum strategy estimates task difficulty by measuring the model’s empirical success rate over 128 attempts per problem, effectively using average success as a proxy for difficulty. While this approach is simple and model-grounded, it is computationally expensive and may not capture the full spectrum of complexity. More sophisticated or efficient alternatives could include model-based difficulty estimation, where a lightweight predictive model learns to estimate difficulty from features such as response length, number of reasoning steps, or uncertainty signals—bypassing the need for repeated sampling. Such approaches could enable faster, more scalable curriculum updates and better generalization across tasks. Future work could explore more adaptive curriculum schedules using richer signals (e.g., loss trends, confidence, or gradient variance), as well as co-adapting the model or optimizer with the curriculum. Finally, understanding when curriculum learning offers diminishing returns, as observed in balanced settings or high-capacity models, is crucial for developing general-purpose and compute-aware curricula.
Appendix A Implementation Details
A.1 Training Configuration
We trained both the actor and critic models using the PPO algorithm. Training was conducted on a single node equipped with 8 A100 GPUs. Each model was trained for approximately 100 steps using the veRL library (Sheng et al., 2024). The key training configuration for Qwen 2.5 7B is shown below. For Qwen 2.5 MATH 1.5B, we set max_response_length=3000 and ulysses_sequence_parallel_size=1 to accommodate its shorter context window of 4096 tokens, while keeping all other parameters unchanged.
export VLLM_ATTENTION_BACKEND=XFORMERS
python3 -m verl.trainer.main_ppo \
data.train_files="$train_files" \
data.val_files="$test_files" \
data.train_batch_size=1024 \
data.max_prompt_length=1024 \
data.max_response_length=8000 \
actor_rollout_ref.model.path=Qwen2.5-7B \
actor_rollout_ref.actor.optim.lr=1e-6 \
actor_rollout_ref.model.use_remove_padding=True \
actor_rollout_ref.actor.ulysses_sequence_parallel_size=2 \
actor_rollout_ref.model.enable_gradient_checkpointing=True \
actor_rollout_ref.actor.ppo_mini_batch_size=1024 \
actor_rollout_ref.actor.use_dynamic_bsz=True \
actor_rollout_ref.actor.ppo_max_token_len_per_gpu=8000 \
actor_rollout_ref.actor.fsdp_config.param_offload=False \
actor_rollout_ref.actor.fsdp_config.optimizer_offload=False \
actor_rollout_ref.rollout.tensor_model_parallel_size=2 \
actor_rollout_ref.rollout.name=vllm \
actor_rollout_ref.rollout.gpu_memory_utilization=0.5 \
actor_rollout_ref.rollout.log_prob_max_token_len_per_gpu=8000 \
actor_rollout_ref.ref.fsdp_config.param_offload=True \
actor_rollout_ref.ref.log_prob_max_token_len_per_gpu=8000 \
critic.optim.lr=1e-5 \
critic.ulysses_sequence_parallel_size=2 \
critic.model.use_remove_padding=True \
critic.model.path=Qwen2.5-7B \
critic.model.enable_gradient_checkpointing=True \
critic.ppo_max_token_len_per_gpu=30000 \
critic.model.fsdp_config.param_offload=False \
critic.model.fsdp_config.optimizer_offload=False \
algorithm.kl_ctrl.kl_coef=0.001 \
trainer.critic_warmup=0 \
trainer.logger=[’console’,’wandb’] \
trainer.project_name=’verl_examples’ \
trainer.experiment_name=’Qwen2.5-7B’ \
trainer.n_gpus_per_node=8 \
+trainer.val_before_train=True \
trainer.nnodes=1 \
trainer.save_freq=10 \
trainer.test_freq=5 \
trainer.total_epochs=15 $@
A.2 AdaRFT Parameters
To enable effective AdaRFT, we tuned several hyperparameters: the training batch size was set to , the target reward to , the sensitivity parameter to , the step size to , and the initial target difficulty to .
The curriculum update mechanism is governed by the difference between the current average reward and the target reward . The difficulty adjustment is computed using a scaled hyperbolic tangent function:
This formulation yields approximately linear updates when , promoting stable learning. When the reward diverges significantly from the target, the function saturates, effectively limiting the magnitude of difficulty updates and preventing destabilizing jumps.
Because the reward is bounded in and the difficulty metric in , we set the step size to ensure alignment in scale. The modulation parameter provides a smooth and controlled progression of difficulty throughout training.