### v2

# Efficient Reinforcement Finetuning via Adaptive Curriculum Learning

Abstract
Reinforcement finetuning (RFT) has shown great potential for enhancing the mathematical reasoning capabilities of large language models (LLMs), but it is often sample- and compute-inefficient, requiring extensive training. In this work, we introduce AdaRFT (Adaptive Curriculum Reinforcement Finetuning), a method that significantly improves both the efficiency and final accuracy of RFT through adaptive curriculum learning. AdaRFT dynamically adjusts the difficulty of training problems based on the model’s recent reward signals, ensuring that the model consistently trains on tasks that are challenging but solvable. This adaptive sampling strategy accelerates learning by maintaining an optimal difficulty range, avoiding wasted computation on problems that are too easy or too hard. AdaRFT requires only a lightweight extension to standard RFT algorithms like Proximal Policy Optimization (PPO), without modifying the reward function or model architecture. Experiments on competition-level math datasets—including AMC, AIME, and IMO-style problems—demonstrate that AdaRFT significantly improves both training efficiency and reasoning performance. We evaluate AdaRFT across multiple data distributions and model sizes, showing that it reduces the number of training steps by up to 2 and improves accuracy by a considerable margin, offering a more scalable and effective RFT framework.
1 Introduction
Reinforcement Finetuning (RFT) has emerged as a powerful technique for aligning large language models (LLMs) with task-specific goals, particularly in domains such as mathematics and code generation where correctness is well defined (DeepSeek-AI et al., 2025; OpenAI et al., 2024b). By optimizing a policy model with reward signals that reflect task success, RFT enables more targeted learning than supervised finetuning (SFT) alone. However, despite its promise, RFT remains sample-inefficient and computationally expensive. Its training involves repeated rollout generation, reward computation, and policy updates—making it costly and difficult to scale (Ahmadian et al., 2024; Kazemnejad et al., 2024; Li et al., 2024; Hu, 2025; Cui et al., 2025). Recent efforts to address RFT inefficiency have focused on algorithmic simplification (e.g., RAFT (Dong et al., 2023), GRPO (DeepSeek-AI et al., 2025), ReMax (Li et al., 2024)), and data-centric strategies (e.g., LIMO (Ye et al., 2025), LIMR (Li et al., 2025)). While these approaches improve sample or compute efficiency, they often introduce trade-offs: algorithmic simplifications may increase variance or limit stability, and static data filtering or scoring can be brittle, computationally heavy, or model-specific. Moreover, most methods’ success relies on fixed datasets or training schedules, which can be suboptimal in non-uniform or imbalanced data regimes. More recently, early efforts have introduced curriculum-like ideas into RFT. Staged curricula divide training into a few manually-defined phases of increasing difficulty (Wen et al., 2025; Luo et al., 2025; Song et al., 2025), but these are coarse-grained and lack adaptivity. Other methods use online data filtering, repeatedly rolling out and pruning training samples until the model’s average reward meets a target threshold (Bae et al., 2025; Yu et al., 2025). While this approach helps prevent the model from stagnating on problems that are either too easy or too difficult, it is not truly adaptive and incurs significant rollout overhead.
To address these limitations, we propose AdaRFT, a reinforcement finetuning method based on adaptive curriculum learning (Bengio et al., 2009), which dynamically adjusts training set difficulty to match the model’s evolving skill level. The intuition is simple: learning is most effective when tasks are neither too easy nor too hard. AdaRFT formalizes this by maintaining a target difficulty level, which increases or decreases based on recent reward feedback. At each step, the model is trained on examples closest to this target, promoting a steady progression through solvable yet challenging tasks. The full algorithm is outlined in Algorithm 1. Unlike prior work that relies on fixed stages, repeated rollouts, or model-specific data processing, AdaRFT is lightweight, general, and model-agnostic. It can be directly applied on top of any standard reinforcement learning (RL) algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017b). We evaluate AdaRFT on a dataset spanning a wide range of competition-level math problems, including AMC, AIME, and IMO-style questions. Across multiple training distributions and two model sizes, AdaRFT significantly improves both training efficiency and final performance. Gains are especially notable in imbalanced data regimes, where static sampling often fails. AdaRFT can reduce training time by up to 2, offering a practical and scalable path to more efficient RFT in structured reasoning tasks.
2 Related Work
Efficient Reinforcement Finetuning.
Most RFT pipelines build on Proximal Policy Optimization (PPO) (Schulman et al., 2017b), with recent variants like RAFT (Dong et al., 2023), ReMax (Li et al., 2024), GRPO (DeepSeek-AI et al., 2025), and REINFORCE++ (Hu, 2025), aiming to reduce computational overhead by simplifying RL components. While effective, these methods often trade off stability or sample efficiency. In parallel, data-centric strategies have emerged as promising alternatives for efficient finetuning. LIMO (Ye et al., 2025) and s1 (Muennighoff et al., 2025) show that small, carefully selected supervised datasets can yield strong downstream performance, but their success hinges on manual curation, prompt engineering, and careful dataset construction, which may not generalize across tasks or models. LIMR (Li et al., 2025) and Wang et al. (2025) proposes scoring training examples based on their estimated learning impact, enabling selective finetuning with fewer samples. Yet, computing these scores requires a full training run, and the scores must be recomputed for each new model, limiting practicality and scalability. In contrast, AdaRFT introduces a lightweight, model-agnostic curriculum learning strategy that dynamically adjusts task difficulty based on reward feedback. This allows continuous adaptation to the model’s capabilities, improving convergence speed and final accuracy without modifying the RL algorithm or requiring manual data curation.
Curriculum Learning for RL.
Curriculum learning (CL) structures training by presenting tasks in an organized progression, typically from easy to hard, to enhance learning efficiency and generalization (Bengio et al., 2009). In RL, CL methods include task sorting by difficulty (Zaremba & Sutskever, 2015; Justesen et al., 2018; Wang et al., 2019), teacher-student frameworks that adaptively select tasks based on learning progress (Matiisen et al., 2017; Portelas et al., 2019), and self-play approaches that induce automatic curricula through agent competition (Sukhbaatar et al., 2018). Other strategies use intermediate-goal generation in sparse-reward settings (Florensa et al., 2018), unsupervised skill discovery (Jabri et al., 2019), or knowledge transfer via progressive networks and imitation (Czarnecki et al., 2018; Rusu et al., 2022). While CL is well-studied in classical RL, its application to RFT of LLMs is still limited. Existing methods typically use staged training with hand-designed difficulty tiers (Wen et al., 2025; Luo et al., 2025; Song et al., 2025), or online filtering schemes that repeatedly sample and discard data until rewards reach a target range (Bae et al., 2025; Yu et al., 2025). These methods either lack adaptability or introduce significant computational overhead due to repeated rollouts. In contrast, AdaRFT is among the first truly adaptive curriculum learning approaches for RFT: it continuously adjusts task difficulty based on the model’s reward signal, enabling efficient, scalable training without fixed schedules or repeated rollouts.
3 AdaRFT
We aim to improve the performance of a policy model for solving mathematical problems through adaptive curriculum learning. Fine-tuning on problems that are too easy or too hard leads to poor learning outcomes. Instead, the model should be trained on problems whose difficulty is close to the model’s current capability. We frame this as an adaptive curriculum learning problem and propose AdaRFT, which adaptively adjusts the target difficulty to keep training problems within a suitable difficulty range. AdaRFT is compatible with a variety of RL algorithms (e.g, GRPO, PPO); in this work, we instantiate it with PPO and refer to this variant as AdaRFT (PPO).
Let be a dataset of mathematical problems, each annotated with a precomputed difficulty score . For example, the MATH dataset (Hendrycks et al., 2021) includes preassigned difficulty levels ranging from 1 to 5. These difficulty scores can also be derived through other methods. One common approach is to use human expert annotations; for instance, problems from the International Mathematical Olympiad (IMO) are generally considered more difficult than those from the American Mathematics Competitions (AMC). Alternatively, difficulty can be estimated based on the model’s empirical success rate: problems that are frequently solved correctly receive lower difficulty scores, while problems that are rarely solved correctly are given higher scores. Another approach is to train a separate model, such as a regression or classification model, to estimate difficulty from features of the problem. Given a predefined difficulty for each problem, the objective is to train a policy that improves its problem-solving ability by dynamically adjusting the training curriculum according to the model’s current performance. Our proposed algorithm, AdaRFT, is shown in Algorithm 1.
3.1 Dynamic Curriculum Sampling
To construct an adaptive curriculum, we define a target difficulty , which represents the current target difficulty level for training. AdaRFT dynamically adjusts based on the model’s reward signal to maintain an optimal difficulty level for learning. At each training step (see Alg. 1), the algorithm computes the absolute difference between the target difficulty and the difficulty of each problem in the dataset (Alg. 1, line 4): for all . The batch of training problems is formed by selecting the problems with the smallest values of (Alg. 1, line 5), producing a batch: . This ensures that the selected problems are closest to the model’s current target difficulty, focusing the learning process on problems that are neither too easy nor too hard.
3.2 Policy Update
The selected batch is used to train the policy model , which generates responses: . A reward signal is computed based on the correctness of the model’s output (Alg. 1, line 7):
3.3 Target Difficulty Update
To adapt the curriculum dynamically, the target difficulty is updated based on the average reward. If the model performs well on the current difficulty level (high reward), the target difficulty increases, making the training problems harder. Conversely, if the model performs poorly, the target difficulty decreases. This dynamic update mechanism lies at the core of AdaRFT’s curriculum adaptation strategy. The update rule (Alg. 1, line 9) is defined as:
Here, are hyperparameters: is the step size for adjusting the target difficulty, controls the sensitivity of the update, and is the target reward level, representing the desired success rate. The function ensures smooth updates and prevents large jumps in difficulty by saturating for large deviations, while the “clip” function constrains the target difficulty within the valid range . These bounds can be manually specified or automatically derived from the training set, for example by taking the minimum and maximum of the difficulty scores . Intuition and guidance for selecting these hyperparameters are discussed in Section 3.4 and 4.3.
3.4 Theoretical Justification for Target Reward
A key component of AdaRFT is its adaptive curriculum mechanism, which steers training toward a target reward level . Intuitively, we aim to train on examples that are neither trivially easy nor prohibitively hard. In this light, setting , corresponding to a success rate of roughly 50%, naturally aligns with this goal. This section formalizes that intuition by analyzing the relationship between reward variance and learnability in RFT with binary rewards.
In entropy-regularized reinforcement learning, the optimal policy can be expressed relative to a reference policy as (Korbak et al., 2022; Go et al., 2023; Rafailov et al., 2023):
where is the inverse temperature parameter controlling entropy regularization, and is the partition function that normalizes the action probability. The corresponding optimal value function and the partition function is given by (Schulman et al., 2017a; Richemond et al., 2024):
We can then take the expectation of the log-ratio between the optimal policy and the initial policy with respect to , leading to (Haarnoja et al., 2017; Schulman et al., 2017a):
Since the left-hand side can be interpreted as the negative reverse KL divergence between and (Rafailov et al., 2024), Bae et al. (2025) show that when the reward with is Bernoulli, the KL divergence is lower-bounded by the reward variance:
where is the model’s success rate on prompt . This implies that the lower bound on the KL divergence, and consequently the gradient magnitude during policy updates, is proportional to the reward variance, which is maximized when . In other words, training on prompts that the model succeeds on roughly half the time may yield the strongest learning signal.
4 Experiments
4.1 Difficulty Estimation
Accurate estimation of problem difficulty is critical for AdaRFT. For difficulty estimation, we use the Qwen 2.5 MATH 7B model (Qwen et al., 2025).111https://huggingface.co/Qwen/Qwen2.5-Math-7B A model that is too strong (e.g., OpenAI o1 (OpenAI et al., 2024b), DeepSeek R1 (DeepSeek-AI et al., 2025)) would solve most problems on the first attempt, leading to poor discrimination between easy and hard problems. Conversely, a model that is too weak (e.g., LLaMA 3.3 1B (Grattafiori et al., 2024)) would fail to solve most problems even after multiple attempts, limiting the signal required for curriculum adaptation.
We select Qwen 2.5 MATH 7B as the estimator because it demonstrates a balanced solving ability. It can solve moderately difficult problems while still failing on more complex problems, thereby providing a more informative difficulty signal. For each problem, the difficulty score is computed as:
where is the number of attempts per problem. In our setup, we use . This approach ensures that difficulty scores reflect the true problem-solving capability of a moderately capable model, preventing over- or under-estimation of difficulty.
To evaluate the stability of our difficulty estimation process, we simulate how confidence varies with different numbers of samples. For each problem, we treat the full set of 128 rollouts as the ground-truth difficulty estimate and compute how often sub-sampled estimates fall within a tolerance of . Specifically, we run 10 random sampling trials per sample size and average the confidence across all problems in the dataset. As shown in Figure 1(a), even with as few as 64 samples, the estimated difficulty remains within of the full estimate over 90% of the time. With just 40 samples, the confidence remains around 80%. These results indicate that accurate and robust difficulty estimation can be achieved with significantly fewer rollouts, reducing the computational burden of large-scale curriculum construction.
To further validate the reliability of our difficulty estimates, we examined their alignment with the difficulty levels provided in the MATH dataset. The MATH dataset comprises 12,500 competition-level mathematics problems sourced from contests such as AMC 10, AMC 12, and AIME. Each problem is categorized into one of five difficulty levels, following the classification system used by the Art of Problem Solving (AoPS) community.222https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki%3ACompetition_ratings In this system, level 1 denotes the easiest problems, while level 5 represents the most difficult. For example, the initial problems in an AMC 8 exam are typically labeled as level 1, whereas AIME problems are often labeled as level 5. As shown in Figure 1(b), there is a clear downward trend in the average solve rate as the labeled difficulty level increases, ranging from 86.0% at level 1 to 52.7% at level 5. Specifically, the AoPS-derived difficulty levels yield a Pearson correlation of () with model success rates. This negative correlation indicates that the model’s empirical performance aligns well with the intended difficulty stratification, reinforcing the utility of both the labeled difficulty levels and our estimation approach in guiding curriculum learning. To further streamline the difficulty estimation process, we also prompted GPT-4o (gpt-4o-0806) (OpenAI et al., 2024a) to assign difficulty levels to the DeepScaleR dataset based on the AoPS rubric. Each problem was presented to GPT-4o with a request to rate its difficulty according to AoPS guidelines (the full prompt is shown in Appendix A.3). This approach provides a lightweight and scalable alternative to rollout-based estimation. As shown in Figure 1(b), GPT-4o’s difficulty ratings also correlate well with the model success rates, with a Pearson correlation of (), making it a practical proxy for curriculum scheduling when computational resources are constrained.
4.2 Dataset
We use the DeepScaleR dataset (Luo et al., 2025) as the training set. DeepScaleR compiles problems from multiple sources, including the American Invitational Mathematics Examination (AIME) from 1984 to 2023 and the American Mathematics Competitions (AMC) prior to 2023. The dataset also includes problems from the Omni-MATH (Gao et al., 2024) and Still datasets (Team, 2025), which feature problems from various national and international math competitions. This results in a diverse and challenging training set, covering a wide range of mathematical domains and difficulty levels.
In practice, we do not have control over the exact difficulty distribution of the data collected for training. This motivates our investigation into how different difficulty distributions influence AdaRFT. To this end, we construct three distinct distributions from the DeepScaleR dataset. The first is a skew-difficult distribution, where most problems are challenging. The second is a skew-easy distribution, where most problems are relatively easy. The third is a uniform distribution, where problems are evenly balanced across all difficulty levels, ensuring a consistent representation of easy, moderate, and hard problems. Each of these three distributions includes 10,000 samples. The data distribution for each setting is shown in Figure 2.
For evaluation, we use six benchmark datasets to assess the model’s performance across different levels of difficulty and mathematical reasoning. The first benchmark, MATH 500 (Lightman et al., 2023), is a subset of the MATH dataset (Hendrycks et al., 2021) containing 500 representative problems designed to test a model’s general mathematical capability. GSM8K (Cobbe et al., 2021) is a set of grade-school math problems. OlympiadBench (He et al., 2024) includes a collection of problems from Olympiad-level mathematics and physics competitions. Minerva Math (Lewkowycz et al., 2022) is a curated set of undergraduate-level math problems that assess complex mathematical reasoning and symbolic manipulation. AMC 23 and AIME 24 include problems from the 2023 American Mathematics Competitions and the 2024 American Invitational Mathematics Examination, respectively. Since AMC 23 contains only 40 problems and AIME 24 only 30, we report accuracy as the average over 8 sampled responses per problem to ensure stable estimates. Together, these datasets span elementary, high school, and advanced competition-level math, providing a comprehensive evaluation of the model’s reasoning abilities.
4.3 Training Setup
We trained two models on the three difficulty-based distributions of the DeepScaleR dataset described in Section 4.2: Qwen 2.5 7B333https://huggingface.co/Qwen/Qwen2.5-7B and Qwen 2.5 MATH 1.5B444https://huggingface.co/Qwen/Qwen2.5-Math-1.5B. This setup allows us to evaluate the effectiveness of AdaRFT on models with different initial performance levels when exposed to skew-hard, skew-easy, and uniform problem distributions.
All models were trained using three different approaches: (1) the standard PPO algorithm, (2) AdaRFT (PPO), our method that integrates adaptive curriculum learning with PPO (see Section 3), and (3) PPO with filtered data, a strong baseline that trains PPO using filtered data based on pass@ accuracy. Following the approach of prior work (Bae et al., 2025; Hu et al., 2025; Zyphra, 2025), we perform a pass@40 analysis for each model and data distribution combination. For the data filtering baseline (3), we remove examples that are either too easy or too hard, excluding problems with solved percentages or . This focuses training on problems of intermediate difficulty. However, this filtering discards a significant portion of the training data, including many potentially useful examples. Moreover, because difficulty is determined via pass@ metrics, filtering must be recomputed each time the model or data distribution changes. In contrast, AdaRFT only requires difficulty estimation once per dataset and adapts to any model during training. For AdaRFT, the target difficulty was dynamically adjusted throughout training based on the model’s reward signal. This adjustment ensured that the model consistently encountered problems that matched its current skill level, preventing the model from being overwhelmed by overly difficult problems or stagnating on problems that were too easy.
The training batch size was set to , with the target reward set to 0.5 to promote learning at a balanced success rate. The sensitivity parameter and step size were tuned using a validation set to ensure stable curriculum updates. We set , , and the initial target difficulty . The step size acts as a scaling factor between the reward signal and the difficulty metric. Since the difficulty metric ranges from 0 to 100 and the reward ranges from 0 to 1, a target reward implies that the maximum reasonable adjustment to the difficulty metric should be around 50. Therefore, we set to scale the reward signal appropriately to the difficulty range. The sensitivity parameter controls the slope of the function. Setting to 2 makes the function behave approximately linearly when the difference between the average reward and the target reward is small. The intuition behind using the function is that when the average reward is close to the target reward, a roughly linear adjustment is appropriate. However, when the average reward deviates significantly from the target reward, linear adjustments may be too large, leading to instability. The function smooths out these adjustments, allowing for more controlled changes when the difference is large while maintaining sensitivity when the difference is small. Both models were trained on 8 A100 GPUs for approximately 100 steps. The implementation details can be found in Appendix A.
5 Result
We evaluate the performance of standard PPO and AdaRFT (PPO) across multiple training setups and two model sizes: Qwen 2.5 MATH 1.5B and Qwen 2.5 7B. Figure 3 presents the learning curves averaged across six benchmarks, while Table 1 and 2 provide a detailed breakdown of accuracy and training efficiency. On average, models trained with AdaRFT (PPO) outperform their PPO-only counterparts in both final accuracy and training efficiency. This improvement is particularly notable in non-uniform data distributions, where curriculum adaptation is most beneficial.
5.1 Training Efficiency
As shown in Figure 3 and Table 1, models trained with AdaRFT consistently require fewer training steps to match the performance of those trained with standard PPO or PPO on filtered data. Specifically, we report how many additional steps are needed for PPO variants to match the performance of AdaRFT at step 60 for Qwen 2.5 Math 1.5B, and step 40 for Qwen 2.5 7B. Because models are evaluated only every 5 training steps, we apply exponential smoothing with a smoothing parameter to the accuracy curves to reduce variance and obtain stable estimates of performance over time. The shaded areas in Figure 3 represent the raw, unsmoothed accuracy 1%, offering a visual cue for the typical fluctuation in evaluation accuracy.
Model | Setup | Method | Avg Step Time (s) | Extra Steps (%) |
Qwen2.5 Math 1.5B | skew-difficult | AdaRFT | 122.24 | +0 (0.0%) |
PPO | 132.95 | +43 (71.7%) | ||
PPO (w/ Filter) | 128.20 | +49 (81.7%) | ||
uniform | AdaRFT | 121.31 | +0 (0.0%) | |
PPO | 126.82 | +34 (56.7%) | ||
PPO (w/ Filter) | 126.35 | +52 (86.7%) | ||
skew-easy | AdaRFT | 120.52 | +0 (0.0%) | |
PPO | 121.15 | +16 (26.7%) | ||
PPO (w/ Filter) | 115.12 | +21 (35.0%) | ||
Qwen2.5 7B | skew-difficult | AdaRFT | 239.92 | +0 (0.0%) |
PPO | 246.21 | +24 (60.0%) | ||
PPO (w/ Filter) | 254.22 | +25 (62.5%) | ||
uniform | AdaRFT | 234.16 | +0 (0.0%) | |
PPO | 243.82 | +13 (32.5%) | ||
PPO (w/ Filter) | 263.11 | +23 (57.5%) | ||
skew-easy | AdaRFT | 247.44 | +0 (0.0%) | |
PPO | 235.27 | +20 (50.0%) | ||
PPO (w/ Filter) | 233.13 | +17 (42.5%) |
For Qwen 2.5 Math 1.5B, standard PPO requires 43 extra steps (+71.7%) in the skew-difficult setting and 34 steps (+56.7%) in the uniform setting to match AdaRFT’s performance. PPO with filtered training data requires even more: +49 steps (81.7%) and +52 steps (86.7%) in the respective settings. In the skew-easy scenario, PPO requires +16 steps (26.7%), while PPO with filtered data needs +21 steps (35.0%) to catch up to AdaRFT. The efficiency gains remain significant with the larger Qwen 2.5 7B model. In the skew-difficult setting, PPO and PPO with filtered data require +24 steps (60.0%) and +25 steps (62.5%), respectively.
In addition to improved sample efficiency, AdaRFT also achieves faster average training time per step across nearly all settings, as reported in Table 1. Though PPO with filtered data can sometimes offer marginal gains in per-step time (e.g., skew-easy setups), it still falls behind in convergence speed. While the underlying causes are discussed in Section 6, these results underscore that AdaRFT is both sample-efficient and compute-efficient, delivering faster and more cost-effective training.
5.2 Model Performance
In addition to improving efficiency, AdaRFT (PPO) also improves the final model performance. As shown in Table 2, at the end of training (step 100), AdaRFT yields consistent improvements in final accuracy across all configurations. The reported averages reflect accuracy across six diverse benchmarks: GSM8K, MATH 500, OlympiadBench, Minerva Math, AMC 23, and AIME 24.
On skew-difficult data, the Qwen 2.5 Math 1.5B model improves from 37.41% (PPO) to 40.48% (AdaRFT (PPO)), a gain of over 3 percentage points in average accuracy. Similar improvements appear in the uniform setting, where AdaRFT (PPO) reaches 41.11%, compared to 37.20% with PPO. Even on skew-easy data, where the baseline performs well, curriculum learning still improves performance, reaching 39.18% versus 38.46%. For the larger Qwen 2.5 7B model, final accuracy gains are also consistent, though slightly more modest. In the skew-difficult setting, AdaRFT (PPO) improves from 44.17% to 46.83%. In the uniform setting, accuracy rises from 44.70% to 46.92%, and in the skew-easy case, from 45.07% to 45.94%. These results show that AdaRFT is effective even for stronger models, enhancing both stability and peak performance.
Model | Setup | Method |
Average | |||||||||||
Qwen 2.5 Math 1.5B | skew-difficult | PPO | 69.67 | 64.60 | 20.65 | 12.87 | 47.50 | 9.17 | 37.41 | |||||||||||
PPO (w/ Filter) | 71.65 | 62.40 | 20.06 | 15.07 | 45.00 | 9.17 | 37.22 | |||||||||||||
AdaRFT | 74.00 | 66.40 | 20.36 | 15.07 | 55.00 | 12.08 | 40.48 | |||||||||||||
uniform | PPO | 71.95 | 65.20 | 21.10 | 15.81 | 42.50 | 6.67 | 37.20 | ||||||||||||
PPO (w/ Filter) | 72.63 | 65.80 | 20.21 | 13.60 | 45.00 | 10.00 | 37.87 | |||||||||||||
AdaRFT | 74.53 | 66.20 | 21.99 | 14.34 | 57.50 | 12.08 | 41.11 | |||||||||||||
skew-easy | PPO | 72.71 | 67.40 | 19.17 | 13.97 | 45.00 | 12.50 | 38.46 | ||||||||||||
PPO (w/ Filter) | 74.75 | 65.20 | 20.36 | 13.60 | 45.00 | 10.00 | 38.15 | |||||||||||||
AdaRFT | 73.62 | 66.20 | 19.91 | 13.97 | 55.00 | 9.17 | 39.18 | |||||||||||||
Qwen 2.5 7B | skew-difficult | PPO | 89.69 | 71.20 | 23.33 | 23.53 | 50.00 | 11.25 | 44.17 | |||||||||||
PPO (w/ Filter) | 88.48 | 72.20 | 24.37 | 25.00 | 50.00 | 12.08 | 45.35 | |||||||||||||
AdaRFT | 90.98 | 71.40 | 25.85 | 22.43 | 52.50 | 15.83 | 46.83 | |||||||||||||
uniform | PPO | 89.31 | 72.40 | 23.63 | 25.37 | 42.50 | 15.00 | 44.70 | ||||||||||||
PPO (w/ Filter) | 89.08 | 74.40 | 23.18 | 22.43 | 45.00 | 13.33 | 44.57 | |||||||||||||
AdaRFT | 90.14 | 72.60 | 24.96 | 24.26 | 55.00 | 14.58 | 46.92 | |||||||||||||
skew-easy | PPO | 89.39 | 73.60 | 23.33 | 24.26 | 47.50 | 13.33 | 45.07 | ||||||||||||
PPO (w/ Filter) | 89.31 | 71.60 | 24.22 | 23.90 | 47.50 | 13.33 | 44.98 | |||||||||||||
AdaRFT | 90.14 | 72.60 | 25.56 | 23.16 | 50.00 | 14.17 | 45.94 |
6 Discussion
6.1 Analysis on Efficiency and Accuracy
Sample efficiency and final performance.
AdaRFT substantially improves both sample efficiency and final model performance, particularly in non-uniform data distributions. By dynamically adjusting task difficulty during training, AdaRFT enables models to allocate learning capacity more effectively, focusing on simpler tasks early on and gradually transitioning to harder ones. This leads to faster convergence and better use of training steps. The performance gains are especially pronounced in the Qwen 2.5 Math 1.5B model, where curriculum structure plays a larger role in guiding learning. For the larger Qwen 2.5 7B model, improvements are still evident, though more modest—likely because the model’s increased capacity allows it to handle diverse and imbalanced tasks without as much external scaffolding.
Reduced training time per step.
AdaRFT also results in a slightly lower average training time per step. This is largely due to the fact that easier problems require fewer tokens to solve. For example, an arithmetic reasoning question from GSM8K might require only around 200 tokens for Qwen 2.5 7B to reach a correct answer, whereas a competition-level math problem from AIME could require around 2000 tokens—a 10 difference in rollout length. The total token length affects multiple components of the training step, including the rollout itself and the subsequent PPO update. While PPO update time does not scale linearly with sequence length due to batching and attention computation patterns, longer sequences still incur higher compute costs. As a result, curriculum learning’s tendency to prioritize shorter, easier problems early in training leads to shorter sequences on average, reducing per-step compute and improving overall training throughput.
When does curriculum learning help?
Our findings show that curriculum learning provides the greatest benefits under two key conditions: (1) imbalanced training distributions, and (2) limited model capacity. In skewed distributions, particularly the skew-difficulty settings, standard PPO often struggles to gain traction early in training due to insufficient reward signals. AdaRFT mitigates this by initially sampling easier problems, enabling the model to bootstrap capabilities before tackling harder content. For example, in the Qwen 2.5 Math 1.5B model trained on skew-difficult data, AdaRFT improves final accuracy by over 3 percentage points and reduces training steps by 44%. This suggests that AdaRFT is especially helpful when the model would otherwise be under-optimized due to overly difficult or unbalanced data.
Conversely, the benefits of AdaRFT are less pronounced when the model is strong enough or the data is already well-balanced. For instance, in the uniform setup with the Qwen 2.5 7B model, AdaRFT offers only modest improvements in convergence speed and final accuracy. Likewise, in the skew-easy setup for the smaller Qwen 2.5 Math 1.5B model, the curriculum yields relatively smaller gains. In both cases, the model is either already exposed to a representative distribution of task difficulties or finds most problems challenging enough, thus reducing the need for dynamic difficulty adjustment.
It is important to note that manual data curation and task scheduling, carefully designed for a specific model, could potentially achieve similar results by selecting a training sequence that aligns with the model’s learning capacity (Yu et al., 2025; Shen et al., 2025; Chen et al., 2025; Zeng et al., 2025). However, such methods require significant human effort, domain knowledge, and often need to be re-tuned for each new model or task. In contrast, AdaRFT requires no manual data curation or model-specific preprocessing. The curriculum automatically adapts to the model’s reward signal during training, making it broadly applicable across model scales and training distributions. This automatic adaptability not only saves engineering effort but also improves scalability and robustness in real-world training pipelines. Moreover, AdaRFT is particularly advantageous in fixed data settings, as it can tailor the difficulty schedule to match the capabilities of any model, whether strong or weak, without altering the dataset, providing a unified solution that generalizes across model architectures and skill levels.
6.2 Data Difficulty on Model Performance
To better understand the effect of data difficulty on model performance, we introduce two additional data distributions: easy-extreme and hard-extreme. Unlike the skew-difficult and skew-easy distributions, which still include a mix of difficulty levels, the easy-extreme and hard-extreme sets consist exclusively of the most polarized examples. Specifically, easy-extreme contains only the easiest samples with difficulty levels no greater than 15, while hard-extreme includes only the hardest samples with difficulty levels of at least 97. Each of these extreme distributions consists of approximately 8,000 samples, providing a focused and controlled evaluation of model behavior under minimal or maximal difficulty conditions. We trained a Qwen 2.5 7B model on each of the two extreme distributions using PPO, and compared their performance to models trained on the uniform distribution with PPO (Uniform) and with AdaRFT instantiated with PPO (Uniform + AdaRFT), as described in Section 5. The results are presented in Figure 4. The key takeaway is that training on only overly easy or hard problems fails to provide useful learning signals, reinforcing the need for AdaRFT to adaptively steer models toward challenges matched to their current ability.
Accuracy.
The leftmost panel of Figure 4 shows that uniform + AdaRFT achieves the highest overall accuracy throughout training, outperforming both uniform and the two extreme settings. This highlights the effectiveness of AdaRFT in guiding the model through an optimal difficulty progression. In contrast, hard-extreme struggles significantly, with a flat and lower trajectory, indicating that exposing the model only to very difficult problems limits learning progress. This suggests that without a gradual exposure strategy, models trained on only the hardest problems are unable to bootstrap their capabilities effectively.
Reward.
The reward trends provide important clues about learning dynamics. The easy-extreme setup achieves the fastest reward improvement during early training, surpassing both uniform and hard-extreme. In particular, easy-extreme consistently operates in a reward range between 0.4 and 0.6 during early training, which corresponds to a success rate that is both challenging and attainable. In contrast, the reward of the uniform and hard-extreme setup lingers below 0.2 in early training, leading to slower learning. This suggests that training on problems with intermediate difficulty—those that are neither trivially easy nor prohibitively hard—provides the most effective learning signal. Notably, AdaRFT is explicitly designed to exploit this insight: by setting the target reward , we encourage the model to train on problems that match this “productive struggle” zone. As shown by the uniform + AdaRFT curve, the algorithm successfully maintains an average reward near 0.5 throughout training, allowing the model to learn at an optimal pace. Notably, while the uniform setup eventually reaches a reward of nearly 0.5 by step 50, it does not result in faster learning. This is likely because the model is already fairly well trained by that stage, so the additional reward signal contributes less to further improvement. In contrast, the hard-extreme model receives almost no reward signal for most of the training, while the uniform setup shows slower and more gradual reward accumulation.
Response Length.
The response length panel reveals how the complexity of generated solutions evolves during training. The hard-extreme model consistently produces the longest responses, with length increasing steadily, reflecting the higher complexity and reasoning depth required by the hardest problems. In contrast, the easy-extreme setup maintains short and stable responses, consistent with its simpler problem set. The uniform and uniform + AdaRFT setups fall between these two extremes. Notably, uniform + AdaRFT shows a gradual increase in response length over time. This trend aligns with the behavior of the curriculum learning algorithm: as the model improves, it is exposed to increasingly difficult problems, which naturally demand more elaborate reasoning and longer solutions. This dynamic suggests that response length can serve as a useful proxy for problem difficulty and reasoning complexity during training.
Difficulty.
Finally, the difficulty panel illustrates how problem difficulty evolves under each setup. The easy-extreme and hard-extreme curves remain flat, confirming that these datasets contain only problems from the tail ends of the difficulty spectrum (i.e., and , respectively). The uniform curve is centered around 50, as expected, while uniform + AdaRFT shows a steady increase in difficulty over time. This adaptive progression confirms that curriculum learning effectively steers the model from easier to harder problems, aligning difficulty with the model’s evolving capabilities.
6.3 Ablation on Target Reward
To better understand the role of the target reward in ADARFT, we perform an ablation study varying in the target difficulty update rule. Recall that controls the target average reward the model is expected to achieve and implicitly steers the curriculum: lower values prioritize easier problems, while higher values shift the curriculum toward more challenging samples. We train a Qwen 2.5 Math 1.5B model on the uniform data distribution with AdaRFT (PPO) using three different values of : 0.2, 0.5, and 0.8. For comparison, we also include standard PPO without ADARFT (denoted as “w/o ADARFT”) as a baseline.
As shown in Figure 5, the model trained with achieves the highest accuracy throughout training. This supports our theoretical motivation in Section 3.4: maximizing reward variance, which occurs when success rate , provides the strongest learning signal. Models with and underperform likely due to curriculum misalignment: overly focuses on easy problems, while overemphasizes difficult ones, both of which limit the model’s capacity to generalize. The reward and difficulty curves align with the accuracy outcomes discussed above. The configuration maintains a stable reward near 0.5, reflecting balanced difficulty exposure. In contrast, results in overly high reward (i.e., easy samples), while maintains a reward around 0.2 for most of training, indicating the model is repeatedly presented with overly difficult problems. As expected, response length is the shortest for and longest for , consistent with the idea that longer responses correlate with problem complexity.
6.4 Training on LLM-Estimated Difficulty
In addition to rollout-based difficulty estimation, we explore an alternative strategy that uses LLM-judged difficulty levels to guide curriculum construction. As described in Section 4.1, we prompt GPT-4o (gpt-4o-0806) to assign difficulty levels to math problems in the DeepScaleR dataset according to the AoPS rubric. This approach offers a lightweight and scalable alternative to computing pass@ success rates from model rollouts, making it especially attractive in low-resource scenarios.
To assess the effectiveness of this strategy, we train a Qwen 2.5 Math 1.5B model on the skew-difficult distribution using AdaRFT (PPO) with two curriculum schedules: one based on rollout-derived pass@ difficulty, and the other guided by GPT-4o’s difficulty ratings. Since the LLM-judged difficulty is on a scale of 1 to 5 (rather than 0 to 100), we set the step size hyperparameter to align the difficulty adjustment magnitude with the reward signal. All other hyperparameters are kept unchanged. As shown in Figure 6, both curriculum strategies outperform standard PPO without curriculum learning. While rollout-based difficulty estimation yields the strongest gains, the LLM-judged curriculum still provides a noticeable improvement over the baseline.
These results demonstrate that AdaRFT remains effective even when the difficulty signal is derived from heuristic or approximate sources like LLM judgments. Although less precise than empirical pass@ metrics, the LLM-based difficulty still provides enough structure to enable meaningful curriculum adaptation. This makes it a practical fallback when rollout computation is too costly, and suggests that future work could explore hybrid approaches that combine lightweight heuristics with periodic empirical calibration.
6.5 AdaRFT with Diverse RL Algorithms
To evaluate the generality of AdaRFT beyond PPO, we trained the Qwen 2.5 Math 1.5B model on a skew-difficult data distribution using two alternative reinforcement learning algorithms: REINFORCE++ and GRPO (see implementation details in Appendix A). As shown in Figure 7, AdaRFT significantly improves both the convergence speed and final accuracy across these variants. Across both cases, the adaptive curriculum acts orthogonally to the underlying optimization method. These results reinforce the plug-and-play nature of AdaRFT: it consistently enhances sample efficiency and policy robustness across algorithmic choices, making it broadly applicable in diverse reinforcement finetuning pipelines. Notably, this generalization holds without any additional tuning or algorithm-specific modifications, underscoring the practical utility of curriculum-aware training in both lightweight and computation-heavy RFT settings.
7 Conclusion
We propose AdaRFT, an adaptive curriculum learning strategy for reinforcement finetuning (RFT) that dynamically matches problem difficulty to a model’s evolving skill level. By adjusting a target difficulty based on reward feedback, AdaRFT improves both sample and compute efficiency without modifying the reward function or underlying RL algorithm. Experiments across multiple data regimes and model sizes show consistent gains in convergence speed and final accuracy, especially in imbalanced training distributions. This lightweight, scalable approach highlights the value of curriculum-aware training for efficient and robust alignment in structured reasoning tasks.
Acknowledgement
The authors would like to thank the members of the LIME lab and the USC NLP group for their helpful insights and feedback.
Appendix A Implementation Details
A.1 Training Configuration
We trained both the actor and critic models using the PPO algorithm on a single node with 8 A100 GPUs. Each model was trained for approximately 100 optimization steps using the veRL library (Sheng et al., 2024). We used two model variants: Qwen2.5-7B and Qwen2.5-MATH-1.5B. The latter has a shorted context window, so we adjusted the max response length and the sequence parallel size accordingly.
Table 3 summarizes the core hyperparameter settings used across all three algorithms: PPO, GRPO, and REINFORCE++. We highlight both shared defaults and algorithm-specific overrides, including KL treatment modes, rollout settings, and critic configurations.
A.2 AdaRFT Parameters
To enable effective AdaRFT, we incorporated a simple curriculum learning mechanism that dynamically adjusts task difficulty during training. We tuned several key hyperparameters: training batch size , target reward , sensitivity parameter , step size , and initial difficulty . These settings were used uniformly across all experiments listed in Table 3. The difficulty is updated as a function of the discrepancy between the running average reward and the target reward , using the following rule:
This formulation promotes stable learning by providing approximately linear updates when , and saturated updates when rewards deviate significantly, thereby avoiding abrupt difficulty shifts. Because the reward is bounded in and the difficulty metric spans , we set the step size to align their scales. The modulation parameter ensures smooth and controlled progression throughout training.
Category | Parameter | PPO | GRPO | REINFORCE++ |
Algorithm-Specific Settings | ||||
General | Advantage estimator | GAE | GRPO | REINFORCE++ |
Gamma () | 1.0 | — | — | |
Lambda () | 1.0 | — | — | |
Batch size | 1024 | 1024 | 1024 | |
Max prompt length | 1024 | 1024 | 1024 | |
Gradient checkpointing | Enabled | Enabled | Enabled | |
Actor | Learning rate | |||
Mini-batch size | 1024 | 1024 | 1024 | |
Dynamic batch size | Enabled | Enabled | Enabled | |
KL penalty role | Reward | Loss | Loss | |
KL loss type | Fixed | Low-variance KL | MSE | |
KL loss coefficient () | 0.001 | 0.001 | 0.001 | |
Entropy coefficient | 0.001 | 0.001 | 0 | |
Clip ratio | 0.2 | 0.2 | 0.2 | |
Gradient clipping | 1.0 | 1.0 | 1.0 | |
Sequence parallel size | Model-specific | Model-specific | Model-specific | |
Rollout | Backend | vLLM | vLLM | vLLM |
Tensor model parallel size | 2 | 2 | 2 | |
Rollouts per sample | 1 | 4 | 1 | |
Nucleus sampling | 1.0 | 1.0 | 1.0 | |
GPU memory utilization | 0.5 | 0.5 | 0.5 | |
Sampling temperature | 1.0 | 1.0 | 1.0 | |
Critic | Warmup steps | 0 | — | — |
Learning rate | — | — | ||
Sequence parallel size | Model-specific | — | — | |
Model-Specific Overrides (shared across all algorithms) | ||||
Qwen2.5-7B | Max response length | 8000 | 8000 | 8000 |
Sequence parallel size | 2 | 2 | 2 | |
Max token length / GPU | 8000 | 8000 | 8000 | |
Qwen2.5-MATH-1.5B | Max response length | 3000 | 3000 | 3000 |
Sequence parallel size | 1 | 1 | 1 | |
Max token length / GPU | 16000 | 16000 | 16000 | |
AdaRFT Parameters | ||||
Curriculum Learning | Target reward () | 0.5 | 0.5 | 0.5 |
Sensitivity () | 2 | 2 | 2 | |
Step size () | 50 | 50 | 50 | |
Initial difficulty () | 0 | 0 | 0 |
A.3 Prompt for Difficulty Estimation Using LLM as a Judge
The prompt used for difficulty estimation (as described in Section 4.1) is shown in Table 4, Table 5, and Table 6. The descriptions of the difficulty scales and examples are sourced from the AoPS Wiki.555https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki:Competition_ratings Although GPT-4o is prompted to rate problem difficulty on a scale from 1 to 10, we found that over 95% of the problems fall within the range of 1 to 5. Therefore, we clip the scores and use a revised scale from 1 to 5. In addition to integer scores, we also allow half-point increments such as 1.5, 2.0, and 2.5 for finer-grained difficulty estimation.